# FLAF: Focal Line and Feature-constrained Active View Planning for Visual Teach and Repeat

### v2

@@ III-B Active View Planning for Feature-based VT&R
The maximum of the cosine functions is achieved when the angle is 0, which indicates that the best view angle is obtained when the focal line of the camera and the mean view angle of the map point overlap.
III-C Path Learning and Tracking with Active Camera
After Teaching, a complete map consisting of plenty of 3D map points and a graph of keyframes are saved with corresponding PTU angles read from the angle encoder. The keyframes with timestamps, poses, and PTU samples contain trajectory information of the active camera and the robot. Then the robot trajectory is computed from PTU samples and camera trajectory by inverse operation of Equation (8):
$\mathbf{X}_{k}=\mathbf{T}_{pt}^{\mathsf{T}}\mathbf{X}_{k}^{\prime}$ (8)
Once the repeat begins, the previously taught feature map is loaded, and the PnP algorithm[22] is executed to determine the initial camera pose. The active camera-based VSLAM, as shown in Fig. 2, is then performed to compute the current pose $\mathbf{X}_{k,w}$. Following this, we search for the closest keyframe pose $X_{r,w}\in\rm{SE}(3)$ ahead of the current localization. Finally, the pose error between $\mathbf{X}_{k,w}$ and $\mathbf{X}_{r,w}$ is processed by a PD controller[23], denoted as $\boldsymbol{C}_{pd}$, to calculate the current velocity $\boldsymbol{\phi}_{k}$:
$\boldsymbol{\phi}{k}=\boldsymbol{C}{pd}(\mathbf{X}{r,w}-\mathbf{X}{k,w})$ (9)
To expedite the reference keyframe search, we define a search window centered on the last reference keyframe, with a fixed width of 10.
@@ III-C Path Learning and Tracking with Active Camera
and 4 are relative and lack a definite scale because the ground truths are derived using SfM with images captured by a monocular camera.
“CR” means the average completion rate in the repeat stage of VT&R. “Time” means the average time used by the sampling-based view planner.
@@
Methods Path1(15.1m) Path2(19.3m) Path3(29.9m) Path4(29.9m)
@@
“SR” indicates the success rate of completing the entire path in the repeat phase.
@@
Figure 5: Trajectories of passive VT&R and active VT&R using different view planning methods are shown. Offsets on the Y-Axis are manually added to separate the overlapping teaching and repeating trajectories for better visualization. The teaching trajectories for Path 1 and Path 2 were obtained via “motion capture” as ground truth, while those for Paths 3 and Path 4 were derived using SfM. Our FLAF-based active VT&R system demonstrates the highest completion rate (CR) across all three paths and can reliably navigate all four paths over multiple loops without failure.
IV Experiments and Discussion
In both phases of VT&R, the active camera operates automatically according to current perception. During the teaching phase, the robot follows a human-guided path and learns the path automatically using the active camera-based VSLAM. Before the repeating, the robot is placed anywhere near the taught path with a similar orientation to the taught one. Once the repeating begins, the robot autonomously runs along the taught path, adjusting the camera orientation as needed.
The main purpose of our experiments is to demonstrate the repeatable and successful VT&R on the challenging paths with our system. As the trajectory errors of repeating are all acceptable, we emphasize the completion and success rate (CR and SR) as the key results, indicating that only our FLAF-based Active VT&R can complete all four paths at a high SR. Notably, the UDVP-based active VT&R consistently fails at the same location across multiple repetitions.
We compared our FLAF-based active view planning against three other methods: Passive VT&R, UDVP-based active VT&R, and FLAF without scoring-based active VT&R. Passive VT&R is achieved on our VT&R system without incorporating active view planning. The UDVP-based active VT&R reproduced the observation model proposed in [16] with our VT&R framework. The comparison with “FLAF without scoring” serves as an ablation study, illustrating the impact of our scoring mechanism as described by equation (7). FLAF without scoring-based active VT&R can also be seen as an active VT&R with the model depicted in [12].
Figure 6: Feature map of path 3 built by our active SLAM and robot views in repeat phase at challenging locations.
IV-A Implementation and Experimental Setup
As shown in Figure 1(d), we fix an Intel Realsense D435 camera on an I-Quotient-Robotics PTU to make up our active camera, which is mounted on a Clearpath-Jackal robot. Our active VT&R system runs in real-time on a notebook with an Intel i7 (2.3GHz) CPU and responds to the images exactly at the frame rate of 20Hz.
Experiments are performed on four paths to evaluate our active view planning method and VT&R system. The first two paths are in the effective range of the motion capture device in the Shenzhen Key Laboratory of Robotics and Computer Vision. Paths 3 and 4 respectively lead the robot from inside the laboratory to a space out of the laboratory and finally back to the start. All the data in Table I are the average results of 10 repeated experiments. The trajectories shown in Fig. 5 and the plots shown in Fig. 7 are a representative selection, considering that the previous methods consistently fail in the same situation. In Fig. 6, we demonstrate the process during which our active VT&R system overcomes the low-texture problem on path 3 that connects different rooms.
Ground truths of paths one and two are obtained from motion capture in the teaching phase. The ground truth of Path 3 is built by SfM, as there is no motion capture outside the laboratory. The visual repeating trajectories of all methods are compared with the ground truths and painted in Fig. 5. In Table I, the CR data shown indicates the VT&R completion rate of this view planning method, and the time data indicates the mean time used for the view planning methods implemented by sampling-based optimization. The efficacy of VSLAM is decreased by the relatively low speed of view planning compared to the SLAM speed of 20-30Hz. The translational RMSE of trajectories is computed using evo[24] by comparing the repeat trajectory with the taught one based on timestamps, resulting in greater error than physical truth. The trajectory consistencies between repeat and teach shown in Fig. 4 are all acceptable for T&R.
On paths 1 and 2, we demonstrate the efficacy of our VT&R system in both an active and a passive way. On paths 3 (Fig. 6) and 4, we show challenging cases with low-texture regions where passive VT&R fails and active VT&R succeeds. Additionally, our FLAF model is verified on all four paths to outperform the existing UDVP in repeating a complete path. FLAF without scoring refers to counting the map points in a range defined by FLAF instead of grading the points by the product of the cosine functions shown in Equation (7).
IV-B Tracking Failure Avoidance Validation
As in Table I and Fig. 5, the passive VT&R achieves a stable and accurate performance on the first two paths but fails in the teaching phase on paths 3 and 4. The few low-texture regions on path one and two are avoided by a considerable human guide. Our active VT&R system with three view planner succeeds in the teaching phase on all 4 paths.
Fig. 6 demonstrates the situations where low-texture regions are overcome by our active VT&R with an active camera that automatically captures informative regions to maintain a stable localization. At position 1, the active camera automatically looks at the poster in the upper left to avoid the white wall. At position 2, the active camera looks up at the ceiling to maintain the localization relying on the square lamps. The robot looks toward the upper right at position 3 to focus on the logo while passing through a low-texture corner. Finally, at position 4, the robot looks toward the upper left at the scroll on the door for abundant features.
@@ IV-C Active View Planning Method Comparison
Active SLAM with different observation models typically succeeds in the teaching phase on all 4 paths. The active camera-based VSLAM has no choice but to see the just mapped points while there are no more feature points triangulated. However, in the repeat phase, existing methods such as UDVP[16] ignore the identifiability of the points by feature detector. Some of the map points (3D coordinates) cannot be identified by VSLAM from a large view angle compared with the view direction in which they were triangulated.
Our FLAF-constrained active view planning outperforms the UDVP method in repeating complete paths because it accounts for the affine change of feature points. The normal line of the map point ($\mathbf{n}_{p}$), shown in Fig. 3, limits the orientation of the active camera in the view angle-invariant range of the feature. As seen in Table I, our method repeats more complete trajectories across all paths, especially in challenging scenarios. Although UDVP achieves more accurate path tracking on path 3, the trajectory errors on all paths by all view planners are negligible for VT&R. Part of the AP-RMSE arises from time-wise point-to-point comparisons, which are challenging to execute precisely. To address this, we modified the timestamps to be uniformly consistent over the same time duration. The trajectories shown in Fig. 5 further confirm that the repeat consistencies with different methods are all pretty good, as the repeated trajectories align well with the taught ones.
@@ V Conclusion
A previous active camera-based VT&R manually steered a gimbal-stabilized camera instead of rotating actively for failure avoidance. Existing active camera-based VSLAM methods did not demonstrate how to integrate their View planners into visual navigation systems. During our tests, the VT&R systems built on these active camera-based VSLAM systems easily failed in the repeat phase without considering the feature-identifiability of the map points. Our proposed focal line and feature (FLAF)-constrained active view planning successfully addressed these failures by considering the view angle difference between the current viewpoint and those at which the map points were triangulated. With our view planner, the active VT&R system successfully finishes all four paths at the highest completion and success rate. Additionally, our point-line plots indicate that the quality of the map points is more important than quantity for stable localization.

### v3

@@ III-C Active View Planning for Feature-based VT&R
The maximum of each cosine function is achieved when the angle is 0, indicating that the best view angle is obtained when the focal line of the camera overlaps with the mean view direction of the map point.
To explain our design, we denote the number of map points within the FoV that can be matched with an ORB feature in $\mathbf{I}_{k}$ as follows:
$N_{S}=f\left(\mathbf{X}_{k},\boldsymbol{M}_{l}\right)$ (6)
which only indicates that $N_{S}$ is a function of the current camera pose and the local map. To increase the number of map points in the central area of the FoV, we designed a scoring function (Eq. 7) based on $\cos({\alpha_{1}})$ (see Fig. 3 for $\alpha_{1}$) to rotate the camera toward regions dense with map points. By adding up the scores of all map points within the FoV, each one of them contributes to the total score of this PTU sample. Map points with smaller $\alpha_{1}$ values receive higher scores, which encourages the clustering of map points in the central area of the FoV. Consequently, $N_{S}$ is increased and tracking failure probability is decreased compared to passive VSLAM. This effect can be represented by the inequality:
$f(\mathbf{T}_{pt}\mathbf{X}_{k})>f(\mathbf{X}_{k})$ (7)
Another goal of our design is to reward the map points with good feature identifiability relative to the current FoV. Feature identifiability refers to the ability of the feature algorithm to recognize the map point from a specific position and orientation. With a complete path map built, there exist more local map points to observe for localization. However, some of the map points can not be identified by the feature detector from an arbitrary position and orientation due to appearance changes[10]. This phenomenon can be summarized as “looking at points visible but not identifiable”, which results in a higher failure rate of repeating.
To address this, we take the view angle ($\alpha_{2}$ in Fig. 3) into consideration and assume the keyframes that observe the map point define a view angle range of successful identification. The mean viewing direction is represented as the feature normal, around which the view angle range is defined. This concept of feature normal was also used in [9, 12]. Mostegel et al.[25] justified using $\cos{\alpha_{2}}$ as the metric of the probability of feature identification to account for the observation of a feature from different viewpoints and view angles. Our scoring function, $\cos({\alpha_{1}})\cdot\cos({\alpha_{2}})$, multiplies two cosine functions to prioritize map points that score highly on both metrics.
III-D Path Learning and Tracking with Active Camera
During teaching, the path map and a robot trajectory are incrementally constructed by VSLAM. After teaching, a complete map consisting of plenty of 3D map points and a graph of keyframes are saved with corresponding PTU angles read from the PTU encoder. The learned trajectory for repeating is stored as a set of key robot poses $\big{\{}\mathbf{X}_{RK}\big{\}}$, each one of which is derived from the keyframe pose and corresponding PTU angles:
$\mathbf{X}_{R,k}=\mathbf{T}_{pt,k}^{-1}\mathbf{X}_{k}$ (8)
Where $\mathbf{X}_{R}$ denotes the robot pose and $\mathbf{X}_{R}$ is the camera pose.
Once repeating begins, the previously taught map is loaded, and the AC-SLAM, as shown in Fig. 5, is performed to localize the robot with the mapping module closed. Meanwhile, the active camera autonomously adjusts its orientation according to real-time perception. The current robot pose during the repeating is computed from the current PTU angles and camera pose by Eq. (8). Following this, we search for the closest robot pose in $\big{\{}\mathbf{X}_{R,i}\big{\}}$ ahead of the current robot pose $\mathbf{X}_{R,k}$ as the current reference robot pose $\mathbf{X}_{r}$. Finally, the pose error between $\mathbf{X}_{r}$ and $\mathbf{X}_{R,k}$ is processed by a PD controller[23], denoted as $\boldsymbol{C}_{pd}$, to calculate the current velocity $\boldsymbol{\phi}_{k}$:
$\boldsymbol{\phi}_{k}=\boldsymbol{C}_{pd}(\mathbf{X}_{r}-\mathbf{X}_{R,k})$ (9)
To expedite the reference keyframe search, a window centered on the last reference keyframe is defined with a fixed width of 10 keyframes.
@@ III-D Path Learning and Tracking with Active Camera
and 4 are relative and lack a definite scale because the ground truths are derived using VSLAM with images captured by a monocular camera.
“CR” means the average completion rate in the repeating stage of VT&R. “Time” means the average time used by the sampling-based view planner.
Figure 6: Trajectories of passive VT&R and active VT&R using different view planning methods are shown. Offsets on the Y-Axis are manually added to separate the overlapping teaching and repeating trajectories for better visualization. The teaching trajectories for Path 1 and Path 2 were obtained via “motion capture” as ground truth, while those for Paths 3 and Path 4 were derived using VSLAM. Our FLAF-based active VT&R system demonstrates the highest completion rate (CR) across all four paths and can reliably navigate all four paths over multiple loops with very few failures.
IV Experiments and Discussion
Our experiments are primarily designed to demonstrate our repeatable and successful VT&R on challenging paths. As the trajectory errors of repeating are all acceptable, we emphasize the completion rate (CR) and success rate (SR), which indicates that only our FLAF-based active VT&R can complete all four paths at a high SR.
IV-A Implementation and Experimental Setup
As shown in Figure 1, we fix an Intel Realsense D435 camera on an I-Quotient-Robotics PTU to make up our active camera, which is mounted on a Clearpath-Jackal robot. Our active VT&R system operates in real-time on a notebook computer equipped with an Intel i7(2.3GHz) processor and responds to the images exactly at the frame rate of 20Hz.
Experiments are performed on 4 paths to evaluate our view planner and VT&R system. The first two paths are in the effective range of the motion capture device in the Shenzhen Key Laboratory of Robotics and Computer Vision. Paths 3 and 4 respectively lead the robot from inside the laboratory to a space out of the laboratory and finally back to the start. All the data in Table I are the average results of 10 repeated experiments. The trajectories shown in Fig. 6, the map shown in Fig. 7, and the plots shown in Fig. 8 are a representative selection, considering that the previous methods consistently fail at a similar location across multiple repetitions.
On Paths 1 and 2, we demonstrate the efficacy of our VT&R system in both an active and a passive way. On Paths 3 (Fig. 7) and 4, we show challenging cases with low-texture regions where passive VT&R fails and active VT&R succeeds. Additionally, our FLAF view planner is verified on all four paths to outperform the existing UDVP in repeating a complete path. FLAF without scoring refers to counting the map points in a range defined by FLAF instead of grading the points by the product of the cosine functions shown in Equation (5).
IV-B Tracking Failure Avoidance Validation
As in Table I and Fig. 6, the passive VT&R achieves a stable and accurate performance on the first two paths but fails in the teaching phase on Paths 3 and 4. The few low-texture regions on Paths 1 and 2 are avoided by a considerable human guide. Our active VT&R system with three view planners succeeds in the teaching phase on all 4 paths.
On Path 3, which connects several rooms, Fig. 7 illustrates how our active VT&R successfully navigates challenging low-texture regions. The active camera autonomously focuses on informative areas, ensuring stable localization throughout. At position 1, the active camera orients toward the poster in the upper left to avoid the white wall. At position 2, the active camera looks up at the ceiling to maintain the localization relying on the square lamps. At position 3, the robot orients toward the upper right to focus on the logo while passing through a low-texture corner. Finally, at position 4, the robot looks toward the upper left at the door for abundant features.
Figure 7: Feature map of Path 3 built by our active SLAM and robot views in repeating phase at challenging locations.
@@ IV-B Tracking Failure Avoidance Validation
Methods Path1(15.1m) Path2(19.3m) Path3(29.9) Path4(19.4)
@@
“SR” indicates the success rate of completing the entire path in repeating.
@@ IV-C Active View Planning Method Comparison
We compared our FLAF-based view planner against three other methods: Passive VT&R, UDVP-based active VT&R, and FLAF without scoring-based active VT&R. The passive VT&R is achieved on our VT&R system without incorporating active view planning. We implement the UDVP-based active VT&R by reproducing the view planner proposed in [16] with our VT&R framework. The comparison with “FLAF without scoring” serves as an ablation study, illustrating the impact of our scoring mechanism as described by equation (5). FLAF without scoring-based active VT&R can also be seen as an active VT&R with the model depicted in [12]. Ground truth of trajectories for Paths 1 and 2 were collected using motion capture during the teaching phase, while for Path 3, VSLAM was used to generate the ground truth, as motion capture was unavailable outside the laboratory.
In Table I, we compare the view planning methods in three metrics: (1) “CR” indicates the VT&R completion rate using different view planning methods, (2) “Time” indicates the average time used for the view planning methods implemented by sampling-based optimization, and (3) “AP-RMSE” indicates the repeating precision by absolute pose-RMSE, which is computed using evo[24] by comparing the repeating trajectory with the taught one based on timestamps, resulting in error data greater than the actual situation. Our FLAF-constrained view planner outperforms the UDVP method in repeating complete paths because it accounts for the affine change of feature points. The normal line of the map point ($\mathbf{n}_{p}$), shown in Fig. 3, limits the orientation of the active camera in the view angle-invariant range of the feature. The efficacy of VSLAM is decreased by the relatively low speed of view planning compared to the SLAM speed of 20Hz.
In Fig. 6, we visualize the trajectories of active VT&R with different view planners, which confirms that the repeated trajectories with different methods all align well with the taught ones. Although the UDVP method achieves more accurate path following on Path 3, the trajectory errors on all paths by all view planners are negligible for VT&R. A portion of the AP-RMSE arises from point-to-point comparisons over time, which are difficult to execute precisely. To address this, we adjusted the timestamps to ensure uniform consistency over the same time duration.
@@ V Conclusion
During our tests, the VT&R systems built on existing view planners frequently failed in the repeat phase without considering the feature-identifiability of the map points. Our proposed focal line and feature (FLAF)-constrained active view planning successfully addressed these failures by considering the view angle difference between the current viewpoint and those at which the map points were triangulated. With our view planner, the active VT&R system successfully finishes all four paths at the highest completion and success rate. Additionally, our point-line plots indicate that the quality of the map points is more important than quantity for stable localization.